OpenStudy (anonymous):

Find f(1), f(2), f(3), and f(4) if f(n) is defined recursively by f(0) = 1 and for n = 0,1,2,... a) f(n+1) = f(n) + 2 b) f(n+1) = 3 f(n)

OpenStudy (gorv):

f(n+1) what value of n we should take so that we get f(1) f(n+1)=f(1)??? at what value on n

OpenStudy (gorv):

what we put in place of n in f(n+1) so that it become f(1)??

OpenStudy (gorv):

yeah you are right

OpenStudy (gorv):

so now f(n+1)=f(n)+1 for n=0 f(1)=f(0)+1 =????

OpenStudy (gorv):

plug in value of f(0)

OpenStudy (anonymous):

f(1) = f(0) + 1 = 1 f(1)?

OpenStudy (gorv):

f(0)=1 f(1)=1+1=2 similarly find others

OpenStudy (anonymous):

The answers in the back of textbook: f (1) = 3, f (2) = 5, f (3) = 7, f (4) = 9 so f(1) is 3 not 2..